OpenStudy (anonymous):

elimination method 5x-2y=13 2x+5y=-18

OpenStudy (across):

I will help you work this one out. Tell me, what is confusing you about it?

hero (hero):

y = (-18 +2x)/5 Set y = y (to eliminate variable y): (13-5x)/-2 = (-18+2x)/5 Cross Multiply: 5(13-5x) = -2(-18+2x) Continue Solving for x: 65 - 25x = 36 -4x 29 = 29x x = 1 Replace x in the original equation to find y: y = -4